# ProbeWalk: Fast Estimation of Biharmonic Distance on Graphs via Probe-Driven Random Walks

### v1

@@ 5. Related Work
The biharmonic distance originated in geometry processing as a Green’s-function-based, spectrum-aware notion of dissimilarity on smooth surfaces (lipman2010Biharmonic). Subsequent work transported the concept to graphs by replacing the continuous bi-Laplacian with the squared graph Laplacian, yielding equivalent formulations via the Laplacian pseudoinverse and truncated spectral expansions (wei2021Biharmonic). This line of study situates biharmonic distance alongside geodesic and resistance distances, while making explicit links to global indices such as the Kirchhoff index and the biharmonic index (klein1993Resistance; wei2021Biharmonic). Recent theory further generalizes biharmonic distance to $k$-harmonic variants and analyzes their role in centrality and clustering on graphs (black2023Understandinga).
It has proved useful well beyond shape analysis. In dynamical networks, biharmonic distance naturally appears in performance and robustness metrics for second-order noisy consensus and related leader-selection objectives (bamieh2012Coherence; fitch2016Joint; yi2022Biharmonic; tyloo2018Robustness). From a learning perspective, the biharmonic distance offers a spectrum-aware metric that emphasizes global structure; together with other spectral distances it underpins positional encodings and spectral clustering, and can be instantiated as a similarity for graph matching—thereby complementing path- and flow-based approaches (kreuzer2021Rethinking; black2023Understandinga; fan2020Spectrala).
@@ Appendix B PROOFS
Step S1: Conditional variance of the U-statistic. Generally, for i.i.d. $U_{1},\dots,U_{R}$ and a symmetric kernel $h:\mathbb{R}^{2}\to\mathbb{R}$, the variance of the second-order U-statistic $U_{R}=\binom{R}{2}^{-1}\sum_{p<q}h(U_{p},U_{q})$ admits the standard decomposition  (hoeffding1948Class)
@@
Let $U=\sum_{g=1}^{G}\mathbf{1}\{|\bar{Y}_{g}-\theta|\geq\varepsilon\theta/4\}$, so that $U\sim\mathrm{Bin}(G,p)$ with $p\leq 1/4$. By Hoeffding’s inequality (one-sided) for Bernoulli sums (hoeffding1963Probability),

### v3

@@ 5. Related Work
The biharmonic distance originated in geometry processing as a Green’s-function-based, spectrum-aware notion of dissimilarity on smooth surfaces (Lipman et al., 2010). Subsequent work transported the concept to graphs by replacing the continuous bi-Laplacian with the squared graph Laplacian, yielding equivalent formulations via the Laplacian pseudoinverse and truncated spectral expansions (Wei et al., 2021). This line of study situates biharmonic distance alongside geodesic and resistance distances, while making explicit links to global indices such as the Kirchhoff index and the biharmonic index (Klein and Randić, 1993; Wei et al., 2021). Recent theory further generalizes biharmonic distance to $k$-harmonic variants and analyzes their role in centrality and clustering on graphs (Black et al., 2023).
It has proved useful well beyond shape analysis. In dynamical networks, biharmonic distance naturally appears in performance and robustness metrics for second-order noisy consensus and related leader-selection objectives (Bamieh et al., 2012; Fitch and Leonard, 2016; Yi et al., 2022; Tyloo et al., 2018). From a learning perspective, the biharmonic distance offers a spectrum-aware metric that emphasizes global structure; together with other spectral distances it underpins positional encodings and spectral clustering, and can be instantiated as a similarity for graph matching—thereby complementing path- and flow-based approaches (Kreuzer et al., 2021; Black et al., 2023; Fan et al., 2020).
@@ Appendix B PROOFS
Step S1: Conditional variance of the U-statistic. Generally, for i.i.d. $U_{1},\dots,U_{R}$ and a symmetric kernel $h:\mathbb{R}^{2}\to\mathbb{R}$, the variance of the second-order U-statistic $U_{R}=\binom{R}{2}^{-1}\sum_{p<q}h(U_{p},U_{q})$ admits the standard decomposition  (Hoeffding, 1948)
@@
Let $U=\sum_{g=1}^{G}\mathbf{1}\{|\bar{Y}_{g}-\theta|\geq\varepsilon\theta/4\}$, so that $U\sim\mathrm{Bin}(G,p)$ with $p\leq 1/4$. By Hoeffding’s inequality (one-sided) for Bernoulli sums (Hoeffding, 1963),